1) The quantity of a substance can be modeled by the function k(t) that satisfies the differential equation  dk/ dt =-1/90 (k-450). One point on this function is k(2) = 900. Based on this model, use a linear approximation to the graph of k(t) at t = 2 to estimate the quantity of the substance at t = 2.1.

 

2) Consider the differential equation dy/dx =x^2y^2with a particular solution y=f(x) having an initial condition y(−3) = 1 . Use the equation of the line tangent to the graph of f at the point (−3, 1) in order to approximate the value of f(−2.8) .

 

3) Given the differential equation dy/dx=-x/y^2 find the particular solution, y=f(x) , with the initial condition f(−6) = −3 .

 

4)  What is the particular solution to the differential equation dy/dx =3 cos(x)y with the initial condition y(pi/2)=-2?




1
Expert's answer
2021-04-29T14:30:40-0400

1) given the differential equation

"\frac {dk}{dt}=\frac {-1}{90} (k-450); k(2)= 900 ...(i)"

is satisfied by thw function k(t) modeling the amount oof the substance; we can use the linear appriximation "k-k(a)= \frac {dk}{dt} (a)(t-a)...(ii)"

centered at [a,k(a)]

in this case, a=2 and k(a)=900.

"\therefore k-900 =\frac{dk}{dt} (2)[t-2]...(iii)"

from (i)"\frac {dk}{dt} (2)=\frac {-1}{90} [900-450]=-5 ...(iv)"

substituting (iv) in (iii)gives;

"k=900-5(t-2) or k=910-5t ...(v)"

as the linear approximation to k(t) satisfying (i)centered at (2, 900).

we can use (v) to approximate k(2.1)as;

"k(2.1) \approx 910-5(2.1)= 899.5"

2) for the differential equation "\frac {dy}{dx} =x^2y^2, y(-3)=1...(i)" where y=f(x)

is the particular solution the tangent line equation at (-3,1) is the linear approximation to f(x) given by;

"y-y(-3)= \frac {dy}{dx} (-3,1)[x+3] or y-1 = \frac {dy}{dx} (-3,1)[x+3] ...(ii)"

from (i), "\frac {dy}{dx} (-3,1)= (-3)^2 (1) =9 ...(iii)"

substituting (iii) in (ii) gives the tangent equation to the curve [y=f(x) satisfying (i)] at (-3,1) to be

"y= 1+9(x+3) or y=9x+28 ...(iv)"

therefore, the approximate value of f(-2.8) is;

"y(-2.8)\approx f(-2.8) \approx 9(-2.8)+28 =2.8"

3) to solve

"\frac {dy}{dx} = \frac x{y^2}, y(-6)=-3 ...(i)"

we first separate the variables to get

"y^2 dy =xdx ...(ii)"

integrating (ii) we have;

"\frac {y^3}3= \frac {x^2}2+c ; y^3 = \frac32 x^2 +3c \implies y=\sqrt[3] {\frac 32 x^2 +k} for k=3c ...(iii)"

given that y(-6)= -3

we have

"-3= \sqrt[3] {\frac 32 (-6)^2 +k} ; -27=54+k \implies k=-81 ...(iv)"

substituting (iv) in (iii) gives the particular solution to (i) as;

"y= \sqrt [3] {\frac 32 x^2 -81}"

4) given; "\frac {dy}{dx} 3cos (x)y; y(\frac {\pi} 2)=-2 ...(i)"

then, separating the variables "\frac {dy}y =3cos (x)dx ...(ii)"

integrating (ii) we have "ln y= 3sinx +c; y= e^{3sinx+c}"

"y= Ae ^{3sinx} for A= e^c ...(iii)"

given "y(\frac {\pi} 2)=-2" then,

"Ae^{3sin \frac {\pi}2}=-2 \implies A=-2e^{-3} ...(iv)"

substituting (iv) in (iii) gives the particular solution to (i) as

"y= -2e^{3sinx-3}"